OpenStudy (anonymous):

simplify 4/5/2/25-5/16

OpenStudy (anonymous):

can you use the draw button and draw the problem can't understand what you have written

OpenStudy (lgbasallote):

is the question \[\huge \frac{\frac 45}{\frac{2}{25}} - \frac{5}{16}\]

OpenStudy (anonymous):

no it is 4/5 then the other two are unter it

OpenStudy (anonymous):